# UUITP-12/25

### v2

@@ 5 Incompleteness in Rindler Wedge
Incompleteness of null geodesics does not necessarily imply the presence of a singularity where the geodesics terminate [12]. In this section, we explore the alternative possibility. Consider the Poincaré patch of AdS3, whose metric is given by
$\mathrm{d}s^{2}=\frac{\mathrm{d}z^{2}-\mathrm{d}x^{+}\mathrm{d}x^{-}}{z^{2}}\,.$ (5.1)
The AdS length scale has been set to one for convenience. We can divide the patch into four Rindler regions depending on the signs of $(x^{+},x^{-})$. We label these regions $\mathcal{R},\mathcal{L},\mathcal{F}$, and $\mathcal{P}$ when the signs are (+, -), (-, +), (+, +), and (-, -) respectively (see [26] for more details on their boundary regions and their importance in the context of AdS/CFT correspondence). We can also cover these Rindler regions using a set of BTZ coordinates $(t,r,\phi)$. The metric in these coordinates takes the form
@@
The map between the BTZ and Poincarè coordinates can be found in [26]. The crucial difference between the BTZ metric and the AdS-Rindler metric in BTZ coordinates (5.2) is that $\phi$ is non-compact in the latter case.
The AdS-Rindler horizon is at $r=1$. The four Rindler regions of the Poincaré patch get mapped to the interior and exterior regions of the BTZ “black hole”. Now, let us look at the growth of extremal surfaces in this coordinate system. We will study surfaces that are symmetric in the $\phi$ coordinate. Since this direction is non-compact, we will regularize the volume by choosing appropriate constant-$\phi$ cutoff surfaces. The metric (5.2) has the same form as (2.1). This allows us to port the uncharged static black hole computations. A trapped extremal surface exists in the interior of the “black hole,” and our singularity theorem tells us that null geodesics are incomplete.
The theorem is signaling the existence of the $r=0$ surface. Clearly, this surface is not a singularity in the full AdS spacetime. Nevertheless, it appears problematic from the perspective of the BTZ coordinate system. To understand the significance of the $r=0$ surface, we must look at the relationship between the Poincaré coordinates and the BTZ coordinates. It turns out that the BTZ coordinates do not cover the entire Rindler region $\mathcal{F}$ [26]. The surface $r=0$ corresponds to the boundary of the BTZ coordinate patch. Beyond this surface, we must use different coordinates to cover the $\mathcal{F}$ patch. Therefore, the geodesic incompleteness stemming from SLC indicates the inability of the coordinate patch to cover the entirety of the Poincaré patch.
@@ 6 Discussion
The growth rate of extremal surfaces has been very useful in signaling the existence of a singularity inside the black hole. A natural question is: what does this object mean in a quantum theory, for instance, in the language of operator growth [29]? Since we are computing the volume of a subregion inside the black hole, a natural guess would be to compute the complexity of a subsystem.
If one can make this notion more precise, we can use these ideas to diagnose the existence of singularities in their holographic duals. Recent developments in the study of operator algebras suggest that the emergence of a Killing horizon in the bulk can be attributed to a particular notion of chaos in the boundary theory [30, 31]. Complexity might provide a natural framework for understanding the emergence of black hole singularities and the interior (see, e.g., [32]). Operator complexity, for example, has been very useful in describing aspects of the exterior geometry of the black hole [33, 34, 35, 36]. Therefore, it is natural to expect complexity to teach us something about the emergence of the black hole interior.

### v3

@@ 5 Incompleteness in Rindler Wedge
Incompleteness of null geodesics does not necessarily imply the presence of a singularity where the geodesics terminate [12]. In this section, we explore the alternative possibility by studying complexity growth in AdS-Rindler wedges. In global coordinates, AdS3 has the metric
$\mathrm{d}s^{2}=\frac{1}{\cos^{2}\rho}\left(-\mathrm{d}\tau^{2}+\mathrm{d}\rho^{2}+\sin^{2}\rho\mathrm{d}\theta^{2}\right)\,.$ (5.1)
We have set the AdS length scale to one for convenience. Using the coordinate transformations,
[…]
we can rewrite the metric in the Poincaré coordinates as follows:
$\mathrm{d}s^{2}=\frac{\mathrm{d}z^{2}-\mathrm{d}x^{+}\mathrm{d}x^{-}}{z^{2}}\,,\qquad\text{where}\ x^{\pm}=t\pm x\,.$ (5.3)
Figure 6: The Poincaré patch of AdS3. The global coordinates cover the entire cylinder. The Poincaré coordinates cover only the solid region shown in the figure, which can be further divided into four Rindler wedges.
The region $z>0$ and $x^{\pm}\in(-\infty,\infty)$ defines the Poincaré patch of the spacetime. We can divide the patch into four Rindler regions depending on the signs of $(x^{+},x^{-})$. We label these regions $\mathcal{R},\mathcal{L},\mathcal{F}$, and $\mathcal{P}$ when the signs are (+, -), (-, +), (+, +), and (-, -) respectively (see Figure 6). The relevance of the boundary of these wedges in the context of AdS/CFT correspondence can be found in [29].
It turns out that we can also cover these Rindler regions using a set of BTZ coordinates $(t,r,\phi)$ given by
$e^{2t}=\frac{x^{+}}{x^{-}},\quad e^{2\phi}=z^{2}-x^{+}x^{-},\quad r^{2}=\frac{z^{2}-x^{+}x^{-}}{z^{2}}\,.$ (5.4)
The metric in these coordinates takes the form
@@
The crucial difference between the BTZ metric and the AdS-Rindler metric in BTZ coordinates (5.5) is that $\phi$ is non-compact in the latter case.
The AdS-Rindler horizon is at $r=1$. The Rindler regions $\mathcal{R},\mathcal{L}$ are mapped to the left and right exterior wedges of the BTZ “black hole.” In contrast, the Rindler regions $\mathcal{F},\mathcal{P}$ are mapped to the future and past interiors, respectively. Now, let us look at the growth of extremal surfaces in this coordinate system. We will study surfaces that are symmetric in the $\phi$ coordinate. Since this direction is non-compact, we will regularize the volume by choosing appropriate constant-$\phi$ cutoff surfaces. The metric (5.5) has the same form as (2.1). This allows us to port the uncharged static black hole computations. A trapped extremal surface exists in the interior of the “black hole,” our singularity theorem tells us that null geodesics are incomplete.
The theorem is signaling the existence of the $r=0$ surface. Clearly, this surface is not a singularity in the full AdS spacetime. Nevertheless, it appears problematic from the perspective of the BTZ coordinate system. To understand the significance of the $r=0$ surface, we must look at the relationship between the Poincaré coordinates and the BTZ coordinates.
Figure 7: (Left) The patch $\mathcal{F}$ is divided into two regions, $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$. The BTZ coordinates cover only $\mathcal{F}_{1}$. (Right) The mapping of various Rindler regions to the corresponding BTZ “black hole.”
It turns out that the BTZ coordinates do not cover the entire Rindler region $\mathcal{F}$ [29]. The BTZ coordinates cover only a part of $\mathcal{F}$, which we will denote by $\mathcal{F}_{1}$ (see Figure 7).
In the region $\mathcal{F}_{2}$, we must use a different set of coordinates (the coordinate transformations can be found in [29]). The surface $r=0$ corresponds to the boundary between $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$. Therefore, the geodesic incompleteness stemming from SLC indicates the inability of the coordinate patch to cover the entirety of the Poincaré patch.
@@ 6 Discussion
The growth rate of extremal surfaces has been very useful in signaling the existence of a singularity inside the black hole. A natural question is: what does this object mean in a quantum theory, for instance, in the language of operator growth [32]? Since we are computing the volume of a subregion inside the black hole, a natural guess would be to compute the complexity of a subsystem. This notion of subsystem complexity should be distinguished from existing ideas in the literature, such as the one in [33], where the volume enclosed by the Ryu-Takayanagi (RT) surface is used to define a notion of subregion complexity in the boundary theory. In contrast, the bulk subregion considered here lies within the black hole interior. It is therefore natural to expect that the associated subregion complexity computes the complexity of boundary operators that are dual to interior degrees of freedom. These operators are inherently non-local in the boundary theory and can, for example, be constructed using the Papadodimas-Raju proposal [34, 35, 36].
If one can make this notion of subregion complexity more precise, we can use these ideas to diagnose the existence of singularities in their holographic duals. Recent developments in the study of operator algebras suggest that the emergence of a Killing horizon in the bulk can be attributed to a particular notion of chaos in the boundary theory [37, 38]. Complexity might provide a natural framework for understanding the emergence of black hole singularities and the interior (see, e.g., [39, 40]). Operator complexity, for example, has been very useful in describing aspects of the exterior geometry of the black hole [41, 42, 43, 44]. Therefore, it is natural to expect complexity to teach us something about the emergence of the black hole interior.